# Escalating LLM-based Code Translation Benchmarking into the Class-level Era

### v2

@@ 5.3. RQ3: Dependency Awareness
The code translation process of LLMs is uncontrollable to some extent, which means it is hard to restrict how LLMs define or name each translated field, variable, and method strictly. Even worse, we also cannot forecast which library LLM will import to implement the equivalent functionality of source PLs. Therefore, it is hard to automatically measure to what extent can LLMs be aware of the necessary dependencies. As such, we conduct human evaluation as a compromise. Specifically, for each translation strategy, we randomly select 20 class-level samples in each LLM’s translation results, a total of 360 selected samples. Afterward, six authors of this paper with over three years of coding experience in Python, Java, and C++ are involved as evaluators and are uniformly divided into three groups. Each group is assigned to one PL-oriented translation for analysis. For the measurement of $d_{F}(\cdot)$ and $d_{M}(\cdot)$ mentioned in Section 4.3, evaluators are required to manually count the number of fields and methods that each caller method relies on. For the measurement of $d_{L}(\cdot)$, we consider three cases as an effective dependency awareness: (1) A translated program invokes the one-to-one-mapping library of the source PL, (2) A translated program invokes one-to-many- or many-to-one-mapping libraries of the source PL, (3) A translated program need not invoke any library, but still implements the equivalent functionality of the source PL. After counting the number of valid dependencies, we compute DEPF, DEPM, and DEPL, respectively, and the corresponding results are shown in Figure 3.
@@
As can be seen, LLMs with the holistic translation strategy perform the best on field dependency awareness, outperforming other strategies by 4.12%–32.80% on average, while LLMs with the min-dependency strategy work much better on invoking useful library dependencies, surpassing others by 7.03%–34.65% on average. As for method dependency awareness, LLMs with holistic and min-dependency strategies perform neck-to-neck, both achieving a DEPM score of around 88% on average. However, LLMs with the standalone translation strategy always perform the worst. A potential explanation is fields need more contextual information to make LLMs observe how different methods exert them to understand their utility, while methods with their implementations to facilitate understanding need relatively less contextual information compared with fields. Therefore, the holistic strategy dominates in DEPF but performs neck-to-neck with the min-dependency strategy in DEPM. As for library dependencies, they are more difficult for LLMs to infer as there are no explicit hints from the source PL, and need a deeper understanding of the program functionalities under translation. With the min-dependency strategy, LLMs can focus more on partial programs, while the holistic strategy brings more noise. In addition, among all LLMs, DeepSeek still performs the best among almost all dependency evaluations and translation strategies, significantly outperforming other LLMs on average by 15.05%, 18.82%, and 20.28% in terms of DEPF, DEPM, and DEPL, respectively.
@@ 5.4. RQ4: Failed Cases Analysis
As the top-performing model among those tested, Deepseek serves as a valuable benchmark for understanding the limitations of state-of-the-art code translation systems. By analyzing its failures, we aim to identify persistent challenges in code translation, highlighting areas where even the most advanced models require further refinement. Therefore, we dedicate 240 person-hours to thoroughly evaluate Deepseek’s 1,397 error cases with the thematic analysis methodology (Cruzes and Dyba, 2011). First, we engage three experts (co-authors in this paper), each with over five years of development experience in Python, Java, and C++. These experts analyze 20% samples of the error cases to identify and summarize the underlying causes. Following this initial analysis, the experts hold a detailed discussion to organize the identified errors into categories, establishing specific criteria and examples for each type. These guidelines are compiled into a codebook, ensuring uniformity and clarity in subsequent evaluations. Subsequently, we recruited eight evaluators with at least one year of experience in Python, Java, and C++. Each sample was assessed through a rigorous, double-blind, dual-review process, where each evaluator independently examined the sample without knowledge of the other’s findings. If both evaluators agreed on the error classification, their result was recorded as final. In cases where evaluations diverged, evaluators engaged in a structured discussion to reach a consensus on the error type and classification. When agreement could not be reached, the case was escalated to the expert team for arbitration, ensuring consistent and accurate categorization. To streamline the evaluation process, we developed a small tool using Python’s tkinter library (Python Documentation, 2024b) to assist evaluators in recording results efficiently. This tool, which we have open-sourced in our repository, allows evaluators to log and categorize errors quickly, reducing manual effort and enhancing overall evaluation efficiency. Throughout the process, if an evaluator identified an error type not covered in the codebook, they were encouraged to report it to the expert team, who would then review and, if necessary, update the codebook to accommodate the new error category. Considering that each sample could contain multiple errors, we categorize on statement/token-level granularity, ensuring the exhaustiveness of the categorization of samples. For ambiguous errors, evaluators discuss to determine the most appropriate classification.
@@
Library-related 12.09 6.06 3.13 2.04 8.04 10.43 11.94 4.50 30.38 17.11 13.82 15.69 14.43 12.17 28.13 17.50 4.29 16.25
Missing required library 10.99 4.55 3.13 2.04 7.14 9.57 11.19 3.60 30.38 17.11 13.16 14.23 14.43 12.17 28.13 17.50 3.07 15.63
API/library mismatch 1.10 1.52 0.00 0.00 0.89 0.87 0.00 0.90 0.00 0.00 0.66 1.46 0.00 0.00 0.00 0.00 1.23 0.00
Copying libraries from source PL 0.00 0.00 0.00 0.00 0.75 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.63 0.00 0.00
Syntax errors 25.27 37.88 1.56 2.04 37.50 42.61 51.49 39.64 5.06 0.00 42.11 30.66 41.24 46.56 3.13 2.50 19.02 33.13
Symbol errors 0.00 0.00 1.56 2.04 13.39 2.61 11.94 9.91 0.00 0.00 12.50 1.82 10.31 8.47 0.00 0.00 0.00 13.44
Language feature errors 16.48 25.76 0.00 0.00 14.29 29.57 25.37 16.22 0.00 0.00 14.47 12.04 12.37 12.70 0.00 0.00 10.43 5.63
Type mismatch errors 8.79 12.12 0.00 0.00 9.82 6.96 2.99 13.51 0.00 0.00 5.92 8.39 3.09 5.29 0.00 0.00 8.59 9.38
Grammar structure errors 0.00 0.00 12.89 13.23 8.21 0.00 0.00 0.00 2.08 2.50 5.06 0.00 0.00 0.87 0.00 4.06 9.21 7.66
Retaining source PL syntax 0.00 0.00 2.58 6.88 2.99 0.00 0.00 0.00 1.04 0.00 0.00 0.00 0.00 2.61 0.00 0.63 0.00 0.73
Function/variable usage issues 35.16 9.09 0.00 4.08 24.11 13.91 23.88 26.13 10.13 0.00 28.29 31.75 29.38 28.57 7.29 0.00 38.65 36.88
Functions missing return values 0.00 0.00 0.00 0.00 2.68 0.00 0.00 0.00 0.00 0.00 0.00 0.73 1.03 0.00 0.00 0.00 5.52 7.19
Return value type mismatch 3.30 0.00 0.00 0.00 3.57 1.74 0.00 0.90 0.00 0.00 2.63 6.20 0.00 1.06 0.00 0.00 4.29 0.31
Attempt to call non-existent function 4.40 1.52 0.00 4.08 2.68 6.09 2.99 1.80 0.00 0.00 7.89 7.66 4.64 3.70 0.00 0.00 17.18 19.06
Attempt to call undefined fields 10.99 3.03 0.00 0.00 12.50 3.48 17.91 11.71 10.13 0.00 17.11 13.14 21.13 22.22 7.29 0.00 9.82 10.31
Parameter errors 16.48 4.55 0.00 0.00 2.68 2.61 2.99 11.71 0.00 0.00 0.66 4.01 2.58 1.59 0.00 0.00 1.84 0.00
Code consistency/completeness 5.49 7.58 0.00 0.00 11.61 6.96 1.49 4.50 0.00 1.32 5.92 17.15 9.79 6.35 0.00 1.25 27.61 5.94
Generated code inconsistency 2.20 3.03 0.00 0.00 7.14 0.87 1.49 4.50 0.00 1.32 4.61 10.22 3.09 2.12 0.00 1.25 7.98 0.63
Non-code content generated 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 3.28 5.67 4.23 0.00 0.00 0.00 1.56
Generated code incomplete 3.30 4.55 0.00 0.00 4.46 6.09 0.00 0.00 0.00 0.00 1.32 3.65 1.03 0.00 0.00 0.00 19.63 3.75
Functional errors 10.99 24.24 95.31 91.84 16.96 7.83 7.46 15.32 54.43 81.58 9.21 3.65 4.64 4.76 61.46 78.75 8.59 3.75
Runtime errors 1.10 3.03 0.00 0.00 0.89 0.00 0.75 0.90 0.00 0.00 0.00 0.00 0.00 0.53 0.00 0.00 0.00 0.00
Other errors 9.89 12.12 0.00 0.00 0.89 18.26 2.99 9.01 0.00 0.00 0.66 1.09 0.52 1.06 0.00 0.00 1.84 4.06
@@
Table 4 presents the evaluation results for three translation strategies, namely holistic, min-dependency, and standalone, across ClassEval-T, categorized by target and source PLs. Major error types include library-related errors, syntax errors, function/variable usage issues, code consistency/completeness errors, generated code functional errors, runtime errors, and Other errors, with each type further divided into specific error types. The results reveal distinct patterns in translation errors based on target and source PL pairings. (1) Functional errors are especially prominent in Python-oriented translations, accounting for 77.23% across all three translation strategies on average. Considering the outstanding performance of DeepSeek in Python-oriented translations mentioned in RQ1, we conclude that other errors in DeepSeek’s Python-oriented translations have been almost eliminated owing to its exceptional capability in code understanding and generation while generating semantic equivalent programs has become its primary challenge. (2) Syntax errors, such as symbol and language feature issues, frequently occur in C++/Java-oriented translations across different strategies, accounting for 40.35% and 34.17%, respectively, which is also revealed by a recent method-level code translation study (Yang et al., 2024a). This pattern can be attributed to the strict syntax and type requirements of these PLs, which pose challenges when translating from more flexible PLs, such as Python. The structural differences between C++ and Java further compound syntax issues when translating between the two. (3) Function and variable usage issues frequently arise in C++/Java-oriented translations across various strategies, constituting 25.37% and 28.93% of the total, respectively. These issues, along with code consistency errors, are especially prevalent in the standalone strategy, where separate translation and missing contextual dependencies limit the LLM’s ability to identify function and field references, increasing the likelihood of calling non-existent functions or fields. Additionally, merging independently translated segments without contextual information often results in logical inconsistencies. To provide a clearer understanding of these types, the following section presents examples illustrating each error type and describes what each entails in practical terms.
@@
Library-related. (1) Missing required library: This category means essential modules or dependencies are omitted during translation. As shown in the example in Figure 4 (a), Java’s in.read() and out.write() are converted to Python without including shutil, which is needed to handle stream copying shutil.copyfileobj(in_stream, out_stream). In RQ3, we found that the min-dependency approach had a high $DEP_{L}$, and here it still results in some missing library errors. This is because, during integration, we include only the translated class skeletons’ libraries and fields, while methods are integrated with their bodies only. Although losing some separately translated dependencies, this solution prevents library conflicts, reduces logical confusion, and simplifies integration. (2) API/library mismatch: This issue occurs when library-specific methods from the source PL are incorrectly applied to the target PL’s libraries. In Figure 4 (b), Java’s contains method in ArrayList is copied directly to C++’s std::list, resulting in an error, as std::list does not support a contains method by default. (3) Copying libraries from source PL: This category includes errors where libraries from the source PL are directly copied into the target PL without proper adaptation. For example (as shown in Figure 4 (c)), Java’s com.fasterxml.jackson.databind.ObjectMapper is directly translated to C++ as #include ¡com/fasterxml/jackson/databind/ObjectMapper.h¿, which does not exist in C++.
@@
Syntax errors. (1) Symbol errors: Symbol errors occur when symbols are missing or misplaced, such as missing parentheses “{” or semicolons “;”, which are required in languages like C++ but not in Python. These missing symbols lead to syntax disruptions when translating between PLs with differing symbol requirements. (2) Language feature errors: These errors arise from misunderstandings of language-specific conventions. For example, as shown in Figure 5 (a), an if condition is incorrectly translated as if (row.find(‘ ’) != npos) instead of the correct statement std::string::npos in C++. (3) Type mismatch errors: Type mismatch errors involve discrepancies in expected types, which lead to incompatible or undefined behavior. Figure 5 (b) presents an example where a HashMap¡Object, Integer¿ is used instead of the expected HashMap¡String, Integer¿, causing type incompatibility in the target PL. (4) Grammar structure errors: Grammar structure errors occur when incorrect syntax structures are used, such as attempting return denominator != 0.0 else 0.0 without an if statement. This structure disrupts the logical flow, as shown in Figure 5 (c). (5) Retaining source PL syntax: This issue occurs when source PL syntax is translated directly into the target PL. For example, as shown in Figure 5 (d), a Python dictionary structure is retained in C++ code, where such structure form is not natively supported.
@@
Function/variable usage issues. (1) Functions missing return values: Errors in this type occur when a function fails to provide a required return value. This leads to incomplete execution of intended logic, causing downstream errors when the function’s output is used in subsequent operations. (2) Return value type mismatch: This error happens when the return type of a function does not align with its expected type. (3) Attempt to call non-existent functions: This issue arises when the target code attempts to call a function that has not been defined or implemented. For example, as shown in Figure 6 (a), the source function search_user_by_username is translated to the target code as searchUserByUsername, but the function searchUserByUsername is not implemented. (4) Attempt to call undefined fields: This error occurs when variables or fields are used in the target code without being defined or declared. For example, as shown in Figure 6 (b), the field books is initialized as a std::vector¡Book¿ and used without prior declaration, causing undefined behavior. (5) Parameter errors: Parameter errors include incorrect parameter types, numbers, or usage in function calls. As shown in Figure 6 (c), an error occurs when the function round is called with two parameters in C++ instead of the required single parameter, resulting in incorrect function usage.
@@
Code consistency/completeness. (1) Generated code inconsistency: Generated code inconsistency occurs when the same data member or function is referenced inconsistently within the same code context, leading to logical errors and a lack of coherence. For example, as shown in Figure 6 (d), the Python function _to_camel_case is translated inconsistently in Java within the same code block: it appears as _to_camel_case in one line and as toCamelCase in another, resulting in compilation errors due to mismatched function names. (2) Non-code content generated: Non-code content refers to instances where irrelevant information, such as “Octal:15” or other unrelated descriptions, is produced instead of code. This also includes cases where the LLM refuses to translate a segment entirely. (3) Generated code incomplete: Generated code incomplete describes translations where, due to token limitations, the code is forcibly truncated before completion, leaving essential parts of the code untranslated.
Functional and Runtime errors. Functional errors occur when the translated code fails to replicate the intended functionality of the source code. These errors stem from misinterpretations in logic, structure, or data handling, resulting in output that deviates from the expected behavior in the target PL. Although the translated code is syntactically correct, functional errors disrupt the code’s ability to perform as intended. Runtime errors are issues that arise during the execution of translated code, leading to unexpected interruptions or failures. These errors prevent the program from running smoothly in the target environment and may result in crashes, incorrect outputs, or halted processes.

### v3

@@ 5.3. RQ3: Dependency Awareness
The code translation process of LLMs is uncontrollable to some extent, which means it is hard to restrict how LLMs define or name each translated field, variable, and method strictly. Even worse, we also cannot forecast which library LLM will import to implement the equivalent functionality of source PLs. Therefore, it is hard to automatically measure to what extent can LLMs be aware of the necessary dependencies. As such, we conduct human evaluation as a compromise. Specifically, for each translation strategy, we randomly select 20 class-level samples in each LLM’s translation results, a total of 480 selected samples. Afterward, six authors of this paper with over three years of coding experience in Python, Java, and C++ are involved as evaluators and are uniformly divided into three groups. Each group is assigned to one PL-oriented translation for analysis. For the measurement of $d_{F}(\cdot)$ and $d_{M}(\cdot)$ mentioned in Section 4.3, evaluators are required to manually count the number of fields and methods that each caller method relies on. For the measurement of $d_{L}(\cdot)$, we consider three cases as an effective dependency awareness: (1) A translated program invokes the one-to-one-mapping library of the source PL, (2) A translated program invokes one-to-many- or many-to-one-mapping libraries of the source PL, (3) A translated program need not invoke any library, but still implements the equivalent functionality of the source PL. After counting the number of valid dependencies, we compute DEPF, DEPM, and DEPL, respectively, and the corresponding results are shown in Figure 3.
@@
As can be seen, LLMs with the holistic translation strategy perform the best on field dependency awareness, outperforming other strategies by 4.00%–24.97% on average, while LLMs with the min-dependency strategy work much better on invoking useful library dependencies, surpassing others by 5.09%–24.04% on average. As for method dependency awareness, LLMs with holistic and min-dependency strategies perform neck-to-neck, both achieving a DEPM score of around 91% on average. However, LLMs with the standalone translation strategy always perform the worst. A potential explanation is fields need more contextual information to make LLMs observe how different methods exert them to understand their utility, while methods with their implementations to facilitate understanding need relatively less contextual information compared with fields. Therefore, the holistic strategy dominates in DEPF but performs neck-to-neck with the min-dependency strategy in DEPM. As for library dependencies, they are more difficult for LLMs to infer as there are no explicit hints from the source PL, and need a deeper understanding of the program functionalities under translation. With the min-dependency strategy, LLMs can focus more on partial programs, while the holistic strategy brings more noise. In addition, among all LLMs, the commercial LLMs still outperform other LLMs in almost all dependency evaluations and translation strategies, surpassing others on average by 16.19%, 15.64%, and 19.58% in terms of DEPF, DEPM, and DEPL, respectively.
@@ 5.4. RQ4: Failed Cases Analysis
As one of the top-performing commercial models, Deepseek serves as a valuable benchmark for understanding the limitations of state-of-the-art code translation systems. By analyzing its failures, we aim to identify persistent challenges in code translation, highlighting areas where even the most advanced models require further refinement. Therefore, we dedicate 240 person-hours to thoroughly evaluate Deepseek’s 1,243 error cases with the thematic analysis methodology (Cruzes and Dyba, 2011), where the error cases are selected from a single-time experiment. First, we engage three experts (co-authors in this paper), each with over five years of development experience in Python, Java, and C++. These experts analyze 20% samples of the error cases to identify and summarize the underlying causes. Following this initial analysis, the experts hold a detailed discussion to organize the identified errors into categories, establishing specific criteria and examples for each type. These guidelines are compiled into a codebook, ensuring uniformity and clarity in subsequent evaluations. Subsequently, we recruit eight evaluators with at least 3 years of experience in Python, Java, and C++. All evaluators are from computer science backgrounds and have achieved scores of no lower than an A in programming courses related to these languages. Each sample is assessed through a rigorous, double-blind, dual-review process, where each evaluator independently examine the sample without knowledge of the other’s findings. If both evaluators agree on the error classification, their result is recorded as final. In cases where evaluations diverge, evaluators engage in a structured discussion to reach a consensus on the error type and classification. When agreement could not be reached, the case is escalated to the expert team for arbitration, ensuring consistent and accurate categorization. To streamline the evaluation process, we develop a small tool using Python’s tkinter library (Python Documentation, 2024b) to assist evaluators in recording results efficiently. This tool, which we have open-sourced in our repository, allows evaluators to log and categorize errors quickly, reducing manual effort and enhancing overall evaluation efficiency. Throughout the process, if an evaluator identify an error type not covered in the codebook, they are encouraged to report it to the expert team, who would then review and, if necessary, update the codebook to accommodate the new error category. Considering that each sample could contain multiple errors, we categorize on statement/token-level granularity, ensuring the exhaustiveness of the categorization of samples. For ambiguous errors, evaluators discuss to determine the most appropriate classification.
@@
Library-related 9.88 6.90 0.00 0.00 12.50 10.71 13.00 9.30 22.22 8.33 13.10 17.27 10.24 7.92 9.78 8.86 20.71 24.19
Missing required library 7.41 5.17 0.00 0.00 8.33 10.71 8.00 6.98 22.22 6.94 13.10 17.27 7.09 5.94 9.78 8.86 17.75 20.97
API/library mismatch 2.47 1.72 0.00 0.00 4.17 0.00 5.00 1.16 0.00 1.39 0.00 0.00 1.57 0.99 0.00 0.00 2.96 3.23
Copying libraries from source PL 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.16 0.00 0.00 0.00 0.00 1.57 0.99 0.00 0.00 0.00 0.00
Syntax errors 30.86 37.93 3.23 0.00 33.33 46.43 45.00 40.70 2.78 4.17 39.29 28.18 43.31 47.52 1.09 1.27 28.40 28.23
Symbol errors 0.00 0.00 0.00 0.00 8.33 7.14 5.00 5.81 0.00 0.00 11.90 6.36 4.72 5.94 0.00 0.00 0.00 8.06
Language feature errors 20.99 24.14 0.00 0.00 4.17 25.00 22.00 20.93 1.39 4.17 13.10 8.18 18.90 24.75 1.09 0.00 21.30 8.87
Type mismatch errors 8.64 13.79 0.00 0.00 8.33 14.29 12.00 13.95 0.00 0.00 9.52 11.82 17.32 14.85 0.00 0.00 3.55 7.26
Grammar structure errors 0.00 0.00 3.23 0.00 8.33 0.00 3.00 0.00 0.00 0.00 3.57 1.82 0.00 0.00 0.00 0.00 2.37 2.42
Retaining source PL syntax 1.23 0.00 0.00 0.00 4.17 0.00 3.00 0.00 1.39 0.00 1.19 0.00 2.36 1.98 0.00 1.27 1.18 1.61
Function/variable usage issues 41.98 17.24 1.61 0.00 25.00 21.43 29.00 27.91 2.78 0.00 28.57 31.82 33.07 30.69 1.09 2.53 36.09 38.71
Functions missing return values 0.00 0.00 0.00 0.00 4.17 3.57 0.00 0.00 0.00 0.00 1.19 0.00 0.00 0.00 0.00 0.00 2.37 2.42
Return value type mismatch 11.11 0.00 0.00 0.00 4.17 7.14 4.00 8.14 0.00 0.00 3.57 6.36 3.15 0.99 0.00 0.00 4.14 3.23
Attempt to call non-existent function 8.64 1.72 0.00 0.00 4.17 7.14 8.00 3.49 0.00 0.00 9.52 10.00 9.45 2.97 0.00 0.00 17.75 20.16
Attempt to call undefined fields 3.70 3.45 1.61 0.00 8.33 0.00 10.00 11.63 2.78 0.00 14.29 13.64 14.17 14.85 1.09 1.27 10.65 12.90
Parameter errors 18.52 12.07 0.00 0.00 4.17 3.57 7.00 4.65 0.00 0.00 0.00 1.82 6.30 11.88 0.00 1.27 1.18 0.00
Code consistency/completeness 3.70 6.90 0.00 0.00 8.33 0.00 4.00 1.16 1.39 0.00 5.95 16.36 3.94 8.91 0.00 1.27 5.33 2.42
Generated code inconsistency 3.70 1.72 0.00 0.00 8.33 0.00 4.00 1.16 1.39 0.00 5.95 13.64 3.94 7.92 0.00 1.27 1.18 0.81
Non-code content generated 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.73 0.00 0.00 0.00 3.28 0.00 0.00 0.00 0.00 0.59 1.61
Generated code incomplete 0.00 5.17 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.99 0.00 0.00 3.55 0.00
Functional errors 6.17 22.41 95.16 100.00 12.50 17.86 2.00 12.79 70.83 87.50 8.33 3.64 4.72 0.99 88.04 86.08 8.28 3.23
Runtime errors 0.00 0.00 0.00 0.00 4.17 0.00 0.00 1.16 0.00 0.00 0.00 0.00 0.00 1.57 0.00 0.00 0.00 0.00
Other errors 7.41 8.62 0.00 0.00 4.17 3.57 7.00 6.98 0.00 0.00 4.76 2.73 3.15 3.96 0.00 0.00 1.18 3.23
@@
Table 4 presents the evaluation results for three translation strategies, namely holistic, min-dependency, and standalone, across ClassEval-T, categorized by target and source PLs. Major error types include library-related errors, syntax errors, function/variable usage issues, code consistency/completeness errors, generated code functional errors, runtime errors, and Other errors, with each type further divided into specific error types. The results reveal distinct patterns in translation errors based on target and source PL pairings. (1) Functional errors are especially prominent in Python-oriented translations, accounting for 87.94% across all three translation strategies on average. Considering the outstanding performance of DeepSeek in Python-oriented translations mentioned in RQ1, we conclude that other errors in DeepSeek’s Python-oriented translations have been almost eliminated owing to its exceptional capability in code understanding and generation while generating semantic equivalent programs has become its primary challenge. (2) Syntax errors, such as symbol and language feature issues, frequently occur in C++/Java-oriented translations across different strategies, accounting for 38.25% and 33.98%, respectively, which is also revealed by a recent method-level code translation study (Yang et al., 2024a). This pattern can be attributed to the strict syntax and type requirements of these PLs, which pose challenges when translating from more flexible PLs, such as Python. The structural differences between C++ and Java further compound syntax issues when translating between the two. (3) Function and variable usage issues, often arising within class structures, frequently occur in C++/Java-oriented translations across various strategies, constituting 33.39% and 30.27% of the total errors, respectively. These issues, along with code consistency errors, are especially prevalent in the standalone strategy, where separate translation and missing contextual dependencies limit the LLM’s ability to identify function and field references, increasing the likelihood of calling non-existent functions or fields. Additionally, merging independently translated segments without contextual information often results in logical inconsistencies. To provide a clearer understanding of these types, the following section presents examples illustrating each error type and describes what each entails in practical terms.
@@
Library-related. (1) Missing required library: This category means essential modules or dependencies are omitted during translation. As shown in the example in Figure 4 (a), the Java code snippet aims to parse a FileReader object to JSONObject one. After the translation to Python code, DeepSeek omits to add the import statement for the corresponding library, i.e., json, but directly uses it, as specified in the red box. In RQ3, we find that the min-dependency approach had a high $DEP_{L}$, and here it still results in some missing library errors. This is because, during integration, we include only the translated class skeletons’ libraries and fields, while methods are integrated with their bodies only. Although losing some separately translated dependencies, this solution prevents library conflicts, reduces logical confusion, and simplifies integration. (2) API/library mismatch: This issue occurs when library-specific methods from the source PL are incorrectly applied to the target PL’s libraries. In Figure 4 (b), a Java method, contains, in WhiteList is copied directly to C++’s std::list, resulting in an error, as std::list does not support the method, contains, by default. (3) Copying libraries from source PL: This category includes errors where libraries from the source PL are directly copied into the target PL without proper adaptation. For example (as shown in Figure 4 (c)), Java’s com.fasterxml.jackson.databind.ObjectMapper is directly translated to C++ as #include ¡com/fasterxml/jackson/databind/ObjectMapper.h¿, which does not exist in C++.
@@
Syntax errors. (1) Symbol errors: Symbol errors occur when symbols are missing or misplaced, such as missing parentheses “{” or semicolons “;”, which are required in languages like C++ but not in Python. These missing symbols lead to syntax disruptions when translating between PLs with differing symbol requirements. (2) Language feature errors: These errors arise from misunderstandings of language-specific conventions. For example, as shown in Figure 5 (a), when implementing Python’s split method in C++, the incorrect statement while (end! = std::npos) is used instead of the correct statement std::string::npos. (3) Type mismatch errors: Type mismatch errors involve discrepancies in expected types, which lead to incompatible or undefined behavior. Figure 5 (b) presents an example where a HashMap¡Object, Integer¿ is used instead of the expected HashMap¡String, Integer¿, causing type incompatibility in the target PL. (4) Grammar structure errors: occur when incorrect syntax structures are used. As shown in Figure 5 (c), the translated Python code attempts to capture an exception with only the code block, except Exception as e:, but ignores the block, try, disrupting the logical flow. (5) Retaining source PL syntax: This issue occurs when source PL syntax is translated directly into the target PL. For example, as shown in Figure 5 (d), Integer is a data structure exclusively used in Java code but is directly copied to C++ code.
@@
Function/variable usage issues. (1) Functions missing return values: Errors in this type occur when a function fails to provide a required return value. This leads to incomplete execution of intended logic, causing downstream errors when the function’s output is used in subsequent operations. (2) Return value type mismatch: This error happens when the return type of a function does not align with its expected type. (3) Attempt to call non-existent functions: This issue arises when the target code attempts to call a function that has not been defined or implemented. For example, as shown in Figure 6 (a), the source function search_user_by_username is translated to the target code as searchUserByUsername, but the function searchUserByUsername is not implemented. (4) Attempt to call undefined fields: This error occurs when variables or fields are used in the target code without being defined or declared. For example, as shown in Figure 6 (b), the field books is initialized as a std::vector¡Book¿ and used without prior declaration, causing undefined behavior. (5) Parameter errors: Parameter errors include incorrect parameter types, numbers, or usage in function calls. As shown in Figure 6 (c), an error occurs when the function palindromic_length is called with two parameters in C++ instead of the required 3 parameters, resulting in incorrect function usage.
@@
Code consistency/completeness. (1) Generated code inconsistency: it occurs when the same variable or function is referenced inconsistently within the same code context, leading to logical errors and a lack of coherence. For example, as shown in Figure 6 (d), the Python function _convert_key is translated inconsistently in Java within the same code block: it appears as _convert_key in one line and as convertKey in another, resulting in compilation errors due to mismatched function names. (2) Non-code content generated: Non-code content refers to instances where irrelevant information, such as “Octal:15” or other unrelated descriptions, is produced instead of code. This also includes cases where the LLM refuses to translate a segment entirely. (3) Generated code incomplete: it describes translations where, due to limitations in model capability or token constraints, essential parts of the code are missing or forcibly truncated before completion, resulting in an incomplete program.
Functional and Runtime errors. Functional errors occur when the translated code fails to replicate the intended functionality of the source code. These errors stem from misinterpretations in logic, structure, or data handling, resulting in outputs that deviate from the expected behavior in the target PL. Although the translated code is syntactically correct, functional errors disrupt the code’s ability to perform as intended. Runtime errors are issues that arise during the execution of translated code, leading to unexpected interruptions or failures. These errors prevent the program from running smoothly in the target environment and may result in crashes, incorrect outputs, or halted processes.